# Understanding Ethical Practices in AI: Insights from a Cross-Role, Cross-Region Survey of AI Development Teams

### v1

@@ 5.4.6. AR + AE Role - Mitigation Strategies
AI researchers develop models, tools, and datasets for research or public use. Therefore, our objective was to understand their practices in mitigating AI risks and applying AI ethics principles.
@@
When asked how often they incorporate AI ethics principles into their research methodology, about 85% mentioned either “always,” “often,” or “sometimes,” while only about 15% said “rarely” or “never.” In terms of seeking guidance on AI ethics in their research, the top three choices are “institutional guidelines and policies” (63.64%), “personal judgment and experience” (59.09%), and “academic literature and journals” (50.0%). Figure 16 shows the result. One participant mentioned “government regulations” as a source of guidance.
@@
As for how they monitor their adherence to ethics statements in their research, the most popular strategies are “regular peer reviews” (59.05%), “feedback from ethics committees” (54.55%), and “internal audits and assessments” (54.55%), while “testing the models’ results” and “reviewing training data for potential biases” appear less significant. This contrasts with the practices of AD and QA roles, where peer reviews and feedback are given lower priority than technical aspects (see Figure 17 for more information).
When asked what challenges they encountered when writing ethics statements for their research papers, 60% said, “difficulty addressing all ethics principles,” and 50% said, “lack of clear guidelines.” These results highlight the need for clearer guidelines and documentation on ethics statements across academic venues that researchers could follow.
Figure 16. AR Roles: Sources of Guidance on AI Ethics in Research
We also asked AI researchers which elements they consider when writing ethics statements in their papers. Among the twelve diverse answers, some had notable considerations, with one participant saying: “Honestly I lie and alway repeat the same 2 sentences that it does not pose any risk whatsoever in different variations. Why nobody is considering the actual risk as the section is an annoyance at best.” Others mentioned being “open about the funding received by the research group,” and “it is usually includes the intended model use case, data collection details (especially if it includes personal data), etc.”
Figure 17. AR Roles: Adherence to Ethics Statements in Research
@@
As shown in Table 1, only three AI ethicists were in group B. We asked them several questions to explore how their role ensures ethics in AI development. When asked how effectively their company implements AI ethics principles in their project, one responded “extremely ineffectively,” while the other two were more positive and mentioned “moderately effectively.” This may suggest that AI ethicists see opportunities for improvement in developing ethical AI technologies. Table 20 shows their company’s actions to raise ethical awareness. The most commonly selected actions are “promote AI ethicists’ roles in the company” and “use case studies to show ethical dilemmas.” Regarding the frequency of evaluating AI systems in their company for potential ethical dilemmas, they only mention either “at a specific project milestone” or “after an incident occurs.” They did not select “continuously,” as part of an ongoing process, or “prior to deploying an AI system.” Yet, a reliable conclusion cannot be drawn due to the small sample size.
Figure 18. AR Roles: Challenges When Writing Ethics Statements
Table 20. AE Roles: Actions Taken by Company to Raise Ethical Awareness
@@
AI researchers commonly incorporate AI ethics principles into their work, drawing on institutional guidelines, personal judgment, and academic literature. Compliance is typically provided through peer review processes and ethics committees. However, many researchers report challenges in fully addressing all ethical dimensions, often citing a lack of precise guidance. AI ethicists also identify room for improvement in how organizations handle AI ethics during development. Nonetheless, they recognize that progress has been made, such as introducing dedicated ethicist roles and using case studies to highlight ethical dilemmas and raise awareness.
5.4.7. Ethical Considerations in AI Training
@@
We asked all participants in Group B several questions about the ethical considerations around training their AI models. 77.43% of them mentioned that they collect, curate, and manage data for training purposes, while the rest were either “Unsure” or said “No.” 84.6% (of those 77.43%) reported that they train their models in-house, and 62.3% (of the 77.43%) mentioned they need “extensive” or “significant resources” to train these models.
@@
Regarding whether their AI systems use personal data for training or testing, 43.81% said “Yes,” while 35.4% and 13.72% said “No” or “Unsure,” respectively. We asked the 101 participants who answered “Yes,” which methods they use to ensure the security and privacy of the data at rest. Figure 19 presents the result. Responses varied, but the top two methods cited were security-related (i.e., “use of encryption” and “access control”), followed by “data masking and anonymization” as the third most-common approach. Policy-related solutions, such as “data retention policies” and “following privacy regulations and standards,” were less common. Interestingly, most participants did not consider backup and recovery protocols (36.36%).
@@
We then followed up with the same group to learn about their practices for ensuring data privacy and preventing leaks. As shown in Figure 20, responses were very diverse; however, “use of encryption” remained the most commonly cited method. “Data masking and minimization” rose from the third to the second most frequently cited method. One interesting observation is the emphasis on “employee training and awareness,” which about half of the participants consider essential. The least popular practice is “third-party vendor management.” These findings highlight an area for improvement, given that third-party code is a major source of data breaches and privacy risks. In contrast, privacy regulations and standards are key in encouraging developers to follow best practices (Prybylo et al., 2024).
@@
Figure 19. Group B - Methods For Ensuring Security and Privacy of Data at Rest
@@
Figure 20. Group B - Methods For Ensuring Private Data Does Not Get Leaked
5.5. Risk and Mitigation Methods Associated with AI Ethics Principles
We asked Group A participants to describe the risks that AI development poses to each of the AI ethics principles within their industry, and Group B participants to describe both these risks and their mitigation strategies for each of the grouped roles.
To categorize their responses, we identified themes and codes following an open-coding process described in Section 3.2. The codebook for a breakdown of the themes can be found in Appendix A - Table 23. The responses revealed some more prevalent codes, including risks due to value alignment of AI systems (e.g., “AI is getting proved daily to be trained in a bias way as long as a human trains it some of their inherent bias will be in the AI for sure”), AI misuse (e.g., “AI could be used to manipulate the facts. This is already seen in AI images spread around for political gain”), and unclear data collection (e.g., “we cannot be sure what data AI systems gather”).
In the following section, we first present Group B participants’ responses for each ethical principle, followed by a summary of Group A’s responses.
5.5.1. Group B - Risk and Mitigation Methods for Each Principle
We asked Group B participants to write about the risks associated with each AI ethics principle in their relevant role.
Regarding the principle Respect for Human Rights, most Group B participants said there are no risks (34.5%) or “Prefer Not to Say” (24.1%). 10.3% of all participants mentioned establishing AI guidelines to mitigate risks, with one participant noting, “Develop and design AI systems in accordance with an AI governance framework.” 5.2% of participants stated risks related to biases and discrimination, with one participant in an AM role saying, “AI models can go rogue and develop their own internal biases.” Another 5.2% said methods related to system monitoring to mitigate risks, with one participant in an AD role emphasizing, “We ensure that our data sets are diverse and representative, avoiding the inclusion of data that could reinforce harmful stereotypes or exclude marginalized groups.”
For the principle Data Protection and the Right to Privacy, 26.5% and 28.6% of Group B participants said there were no risks or “Prefer Not to Say,” respectively. 24.5% mentioned technical security solutions to mitigate risks to this principle, with one participant in an RA role saying, “We ensure data protection and privacy by embedding strict data minimization, encryption, and Technical security solution protocols during requirements gathering.” 11.1% of participants in AD roles emphasized the importance of user consent in mitigating risks associated with this principle.
With regard to the principle Harm Prevention and Beneficence, 29.2% and 27.1% of Group B participants said there were no risks or “Prefer Not to Say,” respectively. 8.3% of responses mentioned ethical guidelines, with some participants in AD roles conducting impact assessments and establishing AI guidelines to mitigate risks associated with this principle. Several participants in ISec roles listed ethical reviews as a way to reduce these risks, with one saying, “A committee is held on a weekly basis to guarantee it.” One participant in an AM role added that “AI failing to deliver promised benefits, leading to missed opportunities for improvement,” while another mentioned the risk of job automation.
Concerning the principle Non-Discrimination and Freedom of Privilege, 29.5% and 27.3% of Group B participants said there were no risks or “Prefer Not to Say,” respectively. 31.8% of participants’ responses were related to biases and discrimination. Several participants in RA roles mentioned the importance of data diversity to mitigate these risks, with one noting: “To ensure non-discrimination and freedom of privileges, we prioritize diverse data representation and rigorously test for bias during requirements gathering and analysis.” Numerous participants from AD roles referred to conducting impact assessments to mitigate these risks, with one responding, “Conduct regular fairness assessments to ensure that the AI system does not discriminate against any group of people.” Several participants highlighted the importance of stakeholder involvement, with two participants in AD roles gathering user feedback.
For the principle Fairness and Justice, 40.8% and 20.4% of Group B participants said there were no risks or “Prefer Not to Say,” respectively. 14.3% of participants’ responses were related to biases and discrimination, with some participants mentioning the importance of data diversity and engaging diverse stakeholders to mitigate these risks. One participant in an AM role responded: “Lack of Diverse Representation: Insufficient diversity in the development team or in data can result in overlooking fairness issues.” Several participants in AD roles stated the importance of stakeholder involvement and gathering user feedback, with one commenting: “We have implemented user feedback mechanisms that allow us to monitor any biases or unfair behaviour that AI could have.”
Regarding the principle Transparency and Explainability, 20.4% and 25.9% of Group B participants said there were no risks or “Prefer Not to Say,” respectively. 11.1% of all participants and 20% of those in AD roles mentioned the importance of open documentation for mitigating risks associated with this principle. 9.3% of all participants and 26.7% of those in AD roles said risk and mitigation methods related to AI design. Some pointed to risks of non-transparent AI (e.g., “Many of the most effective models lack explainability”), while others referred to mitigation methods of using controlled AI models (e.g., “Prioritize AI models that provide explanations for decision making”). 7.4% of all participants and 33.3% of those in RA roles mentioned user feedback to mitigate risks associated with this principle. Several participants from AR roles listed user transparency for mitigating risks related to this principle, with one responding: “We provide thorough explanations for why we built the system and the design choices we made in the process.” Several participants from AD roles stated creating explainability features in AI systems to mitigate these risks, with one adding: “Use of explainable Knowledge Graphs and Knowledge Graph Embeddings at the top level of control.”
Concerning the principle Accountability and Responsibility, 31.4% and 25.9% of Group B participants said there were no risks or “Prefer Not to Say,” respectively. 8.6% of participants’ and 28.6% of those in AM roles’ responses were related to AI design, with one saying: “AI systems are not designed with current regulations in mind, or if they fail to adapt to new regulations, the organization may face legal repercussions. Determining responsibility in such cases can be complex, particularly when AI systems operate.” Others described risks to this principle from non-transparent AI and using controlled AI models to mitigate these risks. 10.7% of participants in AD roles noted regular audits to reduce risks associated with this principle, with one responding: “We implement transparency in our AI systems to ensure the algorithm is auditable.” Another 7.1% of participants in AD roles highlighted the importance of defining roles and responsibilities, with one mentioning: “At the outset, I define clear roles and responsibilities for all team members involved in the AI project.”
In relation to the principle Democracy and Rule of Law, 44.6% and 23.2% of Group B participants said there were no risks or “Prefer Not to Say,” respectively. 22.2% and 14.8% of participants in AD roles mentioned ethical guidelines or following regulations to mitigate risks (e.g., “It is important that intelligent systems should have some railguards when it comes to Human-World-models. Need to encode certain laws regarding democratic solutions”). Several participants’ responses were related to data security and privacy, with one participant in an RA role describing the risk of “AI powered mass surveillance infringing on peoples rights of privacy.” Others listed privacy and other regulations governing AI in their risk mitigation efforts (e.g., “We ensure that our AI systems are compliant with local and international laws, including those governing data protection, privacy, and non-discrimination”).
For the principle Environmental and Social Responsibility, 49% and 19.6% of Group B participants said there were no risks or “Prefer Not to Say,” respectively. The main risk and mitigation methods for this principle relate to resource management (e.g., “Using unnecessary resources that arent sustainable and do not help the environment. Or even using materials that are harmful to the environment”) and energy efficiency (e.g., “Solutions should be environmentally friendly such as not requiring the use of too much energy to run”).
5.5.2. Group A - Risk and Mitigation Methods for Each Principle
We asked Group A participants to write about the risks associated with each AI ethics principle.
For the principle Respect for Human Rights, Group A participants were 24% more likely to describe data security and privacy risks than Group B, with some emphasizing the “risk of over-surveillance.” One participant responded, “AI systems often require large amounts of personal data, raising concerns about data privacy and unauthorized surveillance.” 30.6% of Group A participants indicated a possible value alignment threat (e.g., “AI could not be interested in the wellness of humans”).
Regarding the principle Data Protection and the Right to Privacy, Group A was more likely to describe risks related to big data (e.g., “Since AI uses data and information to operate, it is not certain where data collected by them ends”), data leaks (e.g., “The risks comes in potential data misuse of customer data, unauthorized access to that data, and biased decisions”), and data misuse (e.g., “Companies train AI systems illegally on private data and it’s too hard to control it”).
Concerning the principle Harm Prevention and Beneficence, Group A was more likely to describe risks of AI design and value alignment (e.g., “Capitalist concerns, as usual, take precedence over public well-being”), AI replacing jobs, and overreliance on AI (e.g., “Overreliance on AI systems may lead to humans losing control of key decision-making processes”).
In relation to the principle Non-Discrimination and Freedom of Privileges, Group A was more likely to describe risks related to AI design and value alignment, with several detailing the risks of data sets used in AI development (e.g., “AI is as biased as the data it’s trained on …”).
For the principle Fairness and Justice, Group A participants’ responses were 34.2% and 42.4% more likely to list risks related to biases and discrimination and AI design. Several Group A participants highlighted concerns regarding the value alignment of AI systems, with one responding, “AI-controlled systems can cause physical harm if they malfunction or are designed with flaws.” Others described risks of accessibility to AI systems (e.g., “Inability to access AI hardware”) and the possible misuse and lack of accountability (e.g., “AI is used in fraud detection, it would be unfair to accuse someone of fraud simply because AI suspected it”).
Regarding the principle Transparency and Explainability, Group A participants were 19.5% more likely to emphasize risks associated with black box AI systems and 8.8% more likely to highlight concerns related to unclear data collection practices (e.g., “People and users do not know how data is used and what data is stored”).
With respect to the principle Accountability and Responsibility, Group A participants’ responses were 24% more likely to describe risks related to transparency and explainability, with one participant noting: “The decision-making process of an AI system may involve multiple algorithms and data sources, and determining the responsible party may become complicated when errors or misbehavior occur.”
In relation to the principle Democracy and Rule of Law, Group A participants were 15.3% more likely to associate risks with data security and privacy, with many citing concerns about over-surveillance. They were also more likely to describe risks from AI misuse, with one responding, “AI could be used to manipulate the facts. This is already seen in AI images spread around for political gain.”
For the principle Environmental and Social Responsibility, Group A participants were more likely to list societal risks (e.g., 21.2% are concerned about AI replacing jobs).
Our findings indicate that perceptions of AI ethics risks - and the associated mitigation strategies - differ across principles and roles within AI development teams. Participants identified various risks, including biases, data misuse, ambiguous data collection practices, value misalignment, and potential AI misuse. While some respondents perceived minimal risk for certain principles (e.g., “respect for human rights”), significant concerns were raised regarding fairness, transparency, and accountability. Mitigation strategies were equally varied, encompassing technical measures (such as encryption, bias audits, and curated diverse datasets) and governance practices (including ethical guidelines, regular audits, and stakeholder engagement). These findings align with prior research, underscoring the importance of combining technical and organizational approaches to effectively integrate AI ethics  (Stahl et al., 2022; Kawakami et al., 2024; Agbese et al., 2023). Encouraging communication and establishing clear guidelines throughout the development process, tailored to each role’s responsibilities and context, can support creating more robust and ethically aligned AI systems.
6. Discussion
6.1. Summary of Research Findings
Our findings on the general perception and practices (RQ1) surrounding AI systems indicate that AI is primarily seen as a tool for process automation, performance enhancement, and data synthesis. Participants without experience in AI development tend to be more optimistic about emerging generative AI applications. However, their primary concern centers on ensuring security and privacy in both the use and development of AI tools.
@@ 6.1. Summary of Research Findings
With respect to familiarity (RQ2), we observed considerable variation among AI development teams in their understanding and implementation of ethical principles and governance initiatives. This aligns with prior research emphasizing the importance of involving diverse stakeholders in ethical decision-making processes (e.g., Pant et al. (Pant et al., 2024) and Sanderson et al. (Sanderson et al., 2023)). Our study reinforces the critical role of explainability and transparency in establishing trust in AI systems, echoing findings from Rui et al. (Zhang et al., 2023) and others. Notably, AI managers and requirements analysts exhibit a greater awareness of, and emphasis on, integrating ethical guidelines compared to developers, quality assurance personnel, and information security and privacy teams. Furthermore, individuals working in medium-sized companies, academic institutions, or government agencies tend to report higher familiarity with ethical principles and related initiatives than those employed by small businesses, startups, or large multinational corporations. Geographically, participants from North America, on average, demonstrate greater familiarity than their counterparts in the EU+EEA+UK and other regions. Among all ethical principles, data protection and the right to privacy rank highest in familiarity across roles and locations and are also the most frequently applied in AI development - a finding consistent with the prominence of privacy regulations such as the GDPR (European Union, 2024) and the California Consumer Privacy Act (CCPA) (of California, 2018). Overall, developers maintain a positive outlook on the role of ethics in the future of AI systems. While some expressed concerns that ethics and regulatory frameworks might hinder innovation and competitiveness, particularly for small businesses, the majority remain optimistic about incorporating ethics principles/governance into AI development efforts.
@@
Regarding practice and experience (RQ3), our analysis reveals that roles and demographic factors, such as company size, geographic location, and gender, significantly influence how AI development teams implement various aspects of AI ethics. North American participants are more comfortable with technical solutions and looking towards external support to implement AI ethics, in contrast with those from the EU+EEA+UK are more open to policy-based solutions rooted in regulations and involving diverse stakeholders in decision making. Participants in AM roles demonstrate the strongest commitment to principles related to respecting human rights. In contrast, the relatively limited attention observed among those in QA and ISec roles points to opportunities for strengthening ethical integration in these areas. Notably, AM, QA, and ISec roles are critical to ensuring trust, safety, and security in AI systems and are among the most likely to engage with ethical principles in practice. Participants in AR and AE roles, particularly those working in academic or research institutions, emphasize principles such as accountability and responsibility, fairness and justice, and transparency and explainability. Incorporating their perspectives can help development teams address complex or less visible ethical challenges and contribute to designing AI systems that uphold ethical standards. Finally, our findings underscore the value of gender diversity in AI development teams. Female participants expressed greater concern for ethical principles throughout the AI development process and reported applying them more frequently than their male counterparts.
@@
Concerning risk mitigation strategies (RQ4), we observed various approaches employed by participants across different roles. Overall, those with AI development experience identified key strategies, such as data cleaning to reduce or eliminate bias, ongoing performance monitoring, rigorous testing and validation, and investing in education and training. Role-specific patterns also emerged. Participants in AM roles emphasized ethics training, the implementation of AI ethics guidelines, and collaboration with external experts. Those in AD and QA roles focused on improving training data, reducing bias and discrimination in outputs, and increasing diversity. Their strategies often included evaluating model results, incorporating representative training data, conducting bias audits, and testing outputs for vulnerable groups. ISec professionals, while generally more cautious about AI’s impact on data privacy, still contributed to integrating privacy and security principles into their organizations’ AI development processes. Similarly, ethicists felt their companies could do more to embed and promote AI ethics. Researchers reported regular consideration of ethical concerns, often leveraging peer review and ethics committee feedback, though they also cited challenges, such as unclear guidelines and the tension between ethical goals and research objectives. These findings highlight the need for both technical and organizational strategies to address AI ethics risks and underscore how socio-economic and professional diversity influences the perception and implementation of principles like explainability.
6.2. Research Directions
Building on our findings and the gaps identified in related work, future research should explore interdisciplinary frameworks that bridge technical solutions with organizational processes. We propose four key research directions: (1) Longitudinal studies to examine how enhanced collaboration among diverse roles, particularly management, development, and compliance teams, affects the long-term integration of ethical practices. (2) Development of practical tools and automated frameworks to support “ethics by design” in AI development. (3) Investigation of regulatory and cultural influences, expanding on the work of Rakova et al. (Rakova et al., 2021) and Vakkuri et al. (Vakkuri et al., 2022), to understand how different regulatory environments and organizational cultures shape ethical decision-making. This could inform best practices and enable context-specific interventions. (4) Development of tools and methods for the automated translation of AI ethics regulations into software requirements is essential to address the observed gaps in awareness and compliance. This includes approaches for extracting legal requirements, automating compliance analysis to detect potential violations, and nudging practitioners toward selecting ethical design solutions.
6.3. Educational Takeaway

### v2

@@ 4.4.6 AI Researcher (AR) and AI Ethicist (AE) Roles
AI researchers develop models, tools, and datasets for research or public use. Hence, our objective was to understand their practices in mitigating AI risks and applying AI ethics principles.
@@
When asked how often they incorporate AI ethics principles into their research methodology, $\sim$85% mentioned “always,” “often,” or “sometimes,” while only $\sim$15% said “rarely” or “never.” In terms of seeking guidance on AI ethics in their research, the top three choices are “institutional guidelines and policies” (63.64%), “personal judgment and experience” (59.09%), and “academic literature and journals” (50.0%). One mentioned “government regulations” as a source of guidance (see Figure 15).
@@
As for how they monitor their adherence to ethics statements in their research, the most popular strategies are “regular peer reviews” (59.05%), “feedback from ethics committees” (54.55%), and “internal audits and assessments” (54.55%), while “testing the models’ results” and “reviewing training data for potential biases” appear less significant. This contrasts with the practices of AD and QA roles, where peer reviews and feedback are given lower priority than technical aspects (see Figure 16).
Regarding the challenges encountered when writing ethics statements for their research papers, 60% said, “difficulty addressing all ethics principles,” and 50% said, “lack of clear guidelines.” These results highlight the need for clearer guidelines and documentation on ethics statements across academic venues for researchers to follow.
Figure 15: AR Roles: Sources of Guidance on AI Ethics in Research
AI researchers described which elements they consider when writing ethics statements in their papers. Among the twelve answers, some had notable considerations, with one saying: “Honestly I lie and alway repeat the same 2 sentences that it does not pose any risk whatsoever in different variations. Why nobody is considering the actual risk as the section is an annoyance at best.” Others mentioned being “open about the funding received by the research group,” and “it is usually includes the intended model use case, data collection details (especially if it includes personal data), etc.”
Figure 16: AR Roles: Adherence to Ethics Statements in Research
@@
As shown in Table 1, only three AI ethicists were in Group B. We asked them several questions to explore how their role ensures ethics in AI development. When asked how effectively their company implements AI ethics principles in their project, one responded “extremely ineffectively,” while the other two were more positive, mentioning “moderately effectively.” This may suggest that AI ethicists see opportunities for improvement in developing ethical AI. Table 26 shows their company’s actions to raise ethical awareness, where the most common actions are “promote AI ethicists’ roles in the company” and “use case studies to show ethical dilemmas.” Regarding the frequency of evaluating AI systems for potential ethical dilemmas, they only mention either “at a specific project milestone” or “after an incident occurs.” They did not select “continuously,” as part of an ongoing process, or “prior to deploying an AI system.” Yet, a reliable conclusion cannot be drawn due to the small sample size.
Figure 17: AR Roles: Challenges when Writing Ethics Statements
Table 26: AE Roles: Actions Taken by Company to Raise Ethical Awareness
@@
RQ4 - AR & AE Roles: AI researchers commonly incorporate AI ethics principles into their work, drawing on institutional guidelines, personal judgment, and academic literature. Compliance is typically provided through peer review processes and ethics committees. However, many researchers report challenges in fully addressing all ethical dimensions, often citing a lack of precise guidance. AI ethicists also identify room for improvement in how companies handle AI ethics during development. Nonetheless, they recognize that progress has been made, such as introducing dedicated ethicist roles and using case studies to highlight ethical dilemmas and raise awareness.
4.4.7 Ethical Considerations in Training AI Models
@@
We asked Group B participants questions about the ethical considerations around training their AI models. 77.43% mentioned that they collect, curate, and manage data for training purposes, while the rest were either “Unsure” or said “No.” 84.6% (of those 77.43%) reported that they train their models in-house, and 62.3% (of the 77.43%) mentioned they need “extensive” or “significant resources” to train these models.
@@
Regarding whether their AI systems use personal data for training or testing, 43.81% said “Yes,” while 35.4% and 13.72% said “No” or “Unsure,” respectively. We asked the 101 participants who answered “Yes,” which methods they use to ensure the security and privacy of the data at rest. As Figure 18 shows, the responses varied, but the top two methods cited were security-related (i.e., “use of encryption” and “access control”), followed by “data masking and anonymization” as the third most common approach. Policy-related solutions, such as “data retention policies” and “following privacy regulations and standards,” were less common. Interestingly, most participants did not consider backup and recovery protocols (36.36%).
@@
We then followed up with the same group to learn about their practices for ensuring data privacy and preventing leaks. We received diverse responses, as shown in Figure 19; however, “use of encryption” remained the most commonly cited method. “Data masking and minimization” rose from the third to the second most frequently cited method. One interesting observation is the emphasis on “employee training and awareness,” which about half of the participants consider essential. The least popular practice is “third-party vendor management.” These findings highlight an area for improvement, given that third-party code is a significant source of data breaches and privacy risks. In contrast, privacy regulations and standards are key in encouraging developers to follow best practices [prybylo2024evaluating].
@@
Figure 18: Group B - Methods For Ensuring Security and Privacy of Data at Rest
@@
Figure 19: Group B - Methods For Ensuring Private Data Does Not Get Leaked
5 Discussion
5.1 Summary of Research Findings
Our findings on the general perception of AI (RQ1) surrounding AI systems indicate that AI is primarily seen as a tool for process automation, performance enhancement, and data synthesis. Those without experience in AI development tend to be more optimistic about emerging generative AI applications but are also concerned about ensuring security and privacy in both the use and development of AI tools. Prioritizing privacy aligns with Vakkuri et al. [vakkuri2020aiethics], who found that developers often perceive AI ethics risks related to existing trends in software development (e.g., energy consumption or privacy). On the topic of AI governance initiatives, participants were divided. While the majority saw at least a somewhat positive impact of regulations, a few viewed them as hindrances to progress that tend to favor established corporations. Our results emphasize the importance of AI governance initiatives being better communicated with developers, and for developers to be more proactive in understanding and implementing such initiatives.
@@ 5.1 Summary of Research Findings
We observed considerable variation among AI development teams in their understanding and implementation of ethical principles and governance initiatives (familiarity and predicting factors (RQ2)). This aligns with prior research emphasizing the importance of involving diverse stakeholders in ethical decision-making processes (e.g., Pant et al. [pant2024ethics] and Sanderson et al. [sanderson2023ai]). Our study reinforces the critical role of explainability and transparency in establishing trust in AI systems, echoing findings from Rui et al. [rui2023investigating] and others. Notably, AI managers and requirements analysts exhibit a greater awareness of, and emphasis on, integrating ethical guidelines compared to developers, quality assurance personnel, and information security and privacy teams. Individuals working in medium-sized companies, academic institutions, or government agencies tend to report higher familiarity with principles and initiatives than those employed by small businesses, startups, or large multinational corporations. Among all ethical principles, data protection and the right to privacy rank highest in familiarity across roles and locations, and are applied the most frequently.
LASSO identified familiarity with the EU AI Act (+15.39%) as the strongest predictor of familiarity with AI ethics principles. Mixed-effects models confirmed that while AI development experience significantly predicts familiarity, it accounts for only 6.3% of the variance. In contrast, for how often practitioners consider ethics in their work, familiarity with the AI governance initiatives is the dominant predictor (+0.281), explaining 17.4% of the variance, whereas experience was not selected. LASSO analysis also revealed that positive perceptions of regulation predict both familiarity with principles (+15.23%) and the frequency of ethics consideration (+22.6%). These findings suggest that educating practitioners about AI governance initiatives may effectively enhance both familiarity with and consideration of AI ethics principles, with the EU AI Act showing the strongest association with these benefits. Overall, developers maintain a positive outlook on the role of ethics in the future of AI systems. Although some expressed concerns that ethics and regulatory frameworks could hinder innovation and competitiveness - particularly for small businesses - the majority remain optimistic about integrating ethics principles and governance into AI development practices.
@@
Regarding practice and experience (RQ3), our findings align with Pant et al. [pant2024ethicsai], revealing that roles and demographic factors, such as company size, geographic location, and gender, significantly influence how AI development teams implement various aspects of AI ethics. North American participants are more comfortable with technical solutions and look to external support for implementing AI ethics, whereas those from the EU+EEA+UK are more open to policy-based solutions rooted in regulations and involving diverse stakeholders in decision-making. While these role-based differences exist, random effects analysis revealed that professional role accounts for only 2.9% of variance in knowledge but 15.5% in practice frequency, indicating that organizational and role-based norms shape behavior more significantly than knowledge. Participants in AM roles demonstrate the strongest commitment to principles related to respecting human rights, in contrast to those in QA and ISec roles. Notably, AM, QA, and ISec roles are critical to ensuring trust, safety, and security in AI systems and are among the most likely to engage with ethical principles in practice. Participants in AR and AE roles, particularly those working in academic or research institutions, emphasize principles such as accountability and responsibility, fairness and justice, and transparency and explainability. Incorporating their perspectives can help development teams address complex or less visible ethical challenges and contribute to designing AI systems that uphold ethical standards. Finally, our findings highlight the importance of gender diversity in AI development teams. Female participants expressed greater concern for ethical principles throughout the AI development process and reported applying them more frequently than their male counterparts.
@@
Concerning risk mitigation strategies (RQ4), we observed various approaches employed by participants across different roles. Overall, those with AI development experience identified key strategies, such as data cleaning to reduce or eliminate bias, ongoing performance monitoring, rigorous testing and validation, and investing in education and training. Role-specific patterns also emerged. Participants in AM roles emphasized the importance of ethics training, the implementation of AI ethics guidelines, and collaboration with external experts. Those in AD and QA roles focused on improving training data, reducing bias and discrimination in outputs, and increasing diversity. Their strategies often included evaluating model results, incorporating representative training data, conducting bias audits, and testing outputs for vulnerable groups. ISec professionals, while generally more cautious about AI’s impact on data privacy, still contributed to integrating privacy and security principles into their organizations’ AI development processes. Ethicists felt their companies could do more to embed and promote AI ethics. Researchers reported regularly considering ethical concerns, often leveraging peer review and ethics committee feedback; however, they also cited challenges, such as unclear guidelines and the tension between ethical goals and research objectives. These findings highlight the need for both technical and organizational strategies to address AI ethics risks and underscore how socio-economic and professional diversity influences the perception and implementation of these principles.
5.2 Research Directions
Building on our findings and the gaps identified in related work, future research should explore interdisciplinary frameworks that bridge technical solutions with organizational processes. We propose five key research directions:
(1) Development of standardized threat modeling for ethics in AI systems. Software engineering and privacy research have presented threat modeling frameworks, such as LINDDUN [deng2011privacy], which can capture privacy risks and identify countermeasures early in development. Similar frameworks can be applied to measure and mitigate ethical and regulatory risks in AI systems, enabling developers to understand the risks and threats posed by their applications and providing actionable mitigation methods.
(2) Development of practical tools and automated frameworks to support “ethics by design” in AI development. This addresses Pant et al. [pant2024ethicsai]’s finding that practitioners cite lack of practical tools as a major barrier, and Khan et al. [khan2023aiethics]’s observation that vague principles are difficult to operationalize.
(3) Development of tools and methods for the automated translation of AI ethics regulations into software requirements. This includes approaches for extracting legal requirements, automating compliance analysis to detect potential violations, and encouraging practitioners to select ethical design solutions. As Vakkuri et al. [vakkuri2020aiethics] identified, AI features tend to be treated like regular software features. Translating policy into documentation that aligns with development workflows, such as requirements and user stories [Baldwin2025], can help bridge the gap in developer understanding.
(4) Investigation and development of nudging tools embedded within development environments. Drawing on privacy nudging approaches [balebako2014improving], future research could focus on developing tool-supported methods and approaches that could detect potential AI ethics risks in real-time and prompt consideration of specific mitigation mechanisms.
(5) Tools that enable automated testing of ethics requirements in AI applications. By leveraging recent advancements in LLM-based code analysis, new tools may automatically detect risks, such as biases in training pipelines, explainability in documentation or pull requests, and prompt injection risks in user interfaces. Extending such tools with information retrieval opens the door to analyzing policy alongside applications and providing individualized solutions.
(6) Investigation of regulatory and cultural influences. Building on prior work [rakova2021where, vakkuri2022how], future work could explore how different regulatory environments and organizational cultures shape ethical decision-making. This could inform best practices, such as those outlined by OWASP [owasp2025llm], help create supplementary documentation (similar to that of the EDPB [EDPB_website]), and enable context-specific interventions.
5.3 Educational Takeaways